2nCn * 3^-n = 6C3 * 2^3 * 6^-3
Then n=?
1
Expert's answer
2019-06-11T09:42:16-0400

"_{2n}C_n=_6C_3" so, n=3.

"3^{-n}=\frac{2^3}{6^3}=\frac{1}{3^3}=3^{-3}" so, n=3.

Thus, the only solution is n=3.